OpenStudy (anonymous):

i'm having trouble translating these problems and solving them... "20,000 tickets were sold to the Green Day concert. Stage level seats cost $105 and higher level seats cost $75. If the total money collected from selling tickets was $1,740,000, how many tickets of each type were sold?" "write an equation showing 20,000 tickets were sold." that would be 105x + 75y = 20000 right? then "write an equation showing $1,740,000 in ticket sales" and that's where i start to get confused. then "solve the system of equations to find how many of each type of ticket was sold." thanks!

OpenStudy (melissa_something):

Is this multiple choice?

OpenStudy (anonymous):

no, i have to write the answer myself

OpenStudy (melissa_something):

Ugh I wish I was this good at math. I think you would have to do the same to 1,740,000

OpenStudy (shinalcantara):

let 'x' be the number of tickets for $105 'y' be the number of tickets for $75 since there are 20,000 in total then x + y = 20,000 -----equation 1 total money collected is $1,740,000 then you'll have the equation: 105x + 75y = 1,740,000 ------equation 2 from equation 1 x + y =20,000 x = 20,000 - y ----equation 1' substitute equation 1' to equation 2 105x + 75y = 1,740,000 105(20,000 - y) + 75y = 1,740,000 2,100,000 - 105y + 75y = 1,740,000 -30y = -360,000 y = 12,000 substitute y=12,000 to equation 1' x = 20,000 - y x = 20,000 - 12,000 x = 8,000